Many counties in Florida missed many school days in the fall of 2004 due to hurricanes that year. A solution for how to make up those days was to add time to each school day for a portion of the year. It was decided that in order to make up a 6 hour school day, 10 minutes would be added to each day. How many extended days would be needed to make up 3 school days?
Answer by stanbon(75887)

Many counties in Florida missed many school days in the fall of 2004 due to hurricanes that year. A solution for how to make up those days was to add time to each school day for a portion of the year. It was decided that in order to make up a 6 hour school day, 10 minutes would be added to each day. How many extended days would be needed to make up 3 school days?
6 hrs = 6*60 = 360 minutes
--------
3 school days = 3*360 = 1080 min
----------------
# of days = 1080 min/10 min = 108 days
---------------
Cheers,
Stan H.
-----------